OpenStudy (anonymous):

help me pwease =( Women in the United States continue to be paid less than men overall. Nationally, in 2007, the median weekly earnings of female full-time wage and salary workers totaled $614, while it was $766 for males. In 2007, male full-time wage and salary workers earned _________ % more than female full-time wage and salary workers. In 2007, female full-time wage and salary workers earned _________ % less than male full-time wage and salary workers.

OpenStudy (anonymous):

i want a second opinion

OpenStudy (anonymous):

y thank u =D

OpenStudy (anonymous):

haha

OpenStudy (saifoo.khan):

Are we playing a tag game?

OpenStudy (anonymous):

can u help me out with this problem please =D

OpenStudy (unklerhaukus):

766/614-1=

OpenStudy (anonymous):

is this correct? Males earned 766/1380= 55.51% more. Females made 614/1380= 44.49% less.

OpenStudy (unklerhaukus):

no

OpenStudy (aripotta):

i got that males earned about 25% more

OpenStudy (anonymous):

766−614/614 = 24.76%

OpenStudy (unklerhaukus):

betta

OpenStudy (aripotta):

omg my eyebrow itches

OpenStudy (anonymous):

if thats the case then what percent of male full-time wage and salary worker median weekly earnings does the female median weekly earnings represent?

OpenStudy (aripotta):

lol

OpenStudy (anonymous):

what percent of male full-time wage and salary worker median weekly earnings does the female median weekly earnings represent?